# Sparse Bayesian multidimensional scaling(s)

### v1

@@ 2.3 Posterior consistency
$\delta_{nn^{\prime}}\sim N_{(0,M)}(\delta_{nn^{\prime}}^{*},\sigma^{2}),\ n\neq n^{\prime},\ n,n^{\prime}=[N],$ (6)
a generalization of (1) insofar as M can be any number within the interval $(0,\infty)$. The posterior density function of the unknown parameters ($\mathbf{X},I,\sigma^{2},M$) is proportional to $\mathcal{L}(\bm{\Delta}|\mathbf{X},\sigma^{2},M)$, the BMDS likelihood function of model (6), and the priors put on each auxiliary parameter, e.g.,
@@
We examine the posterior consistency of subsampled distance matrices under simple conditions. Fixing some interval $I$, we sample points $x_{1},\dots,x_{N}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N_{I}(0,1)$. Let $\Delta^{*}$ be the associated distance matrix with entries $\delta^{*}_{nn^{\prime}}=|x_{n}-x_{n^{\prime}}|$ and $\delta_{nn^{\prime}}$ be the truncated noisy observations of this matrix sampled from model (6). We set a prior on $I$, $M$ and $\sigma^{2}$ that has compact support and is bounded away from 0 and infinity on its support. In addition, we fix in advance a collection of indices $J_{n,N}\subset[N]\backslash\{n\}$ of observations to keep for each object $n$, treating this choice as non-random in the following. Next, we make some assumptions about which observations are kept.
@@
Fix $K\in\mathbb{N}$. Assume there exists a sequence $\{\ell_{N}\}_{N\in\mathbb{N}}$ and a collection of partitions $\{G_{n,N}^{(k)}\}_{k=1}^{K}$ of $J_{n,N}$ with the following properties:
@@
We verify that Assumption 1 holds for L-sBMDS given $L\geq(\ell_{n}+1)K$ landmarks and $L\ll N$. We can think of $\ell_{n}$ as the number of retained entries in the sparsest row (up to a universal constant). Let $\ell_{N}=ceiling(\frac{\sqrt{N}}{2})$, so that it satisfies part 3 of Assumption 1. For all $n\in[N],|J_{n,N}|\geq L$ and for objects $n,n^{\prime}\in[N],|J_{n,N}\cap J_{n^{\prime},N}|\geq L-1$. We can then divide $J_{n,N}$ into a collection of $K$ partitions such that each partition is approximately of size $\frac{L}{K}$, satisfying part 1 of Assumption 1. For example, let $G_{n,N}^{(k)}=\{L^{*}\}_{L^{*}=floor(\frac{(k-1)L}{K})+1}^{floor(\frac{kL}{K})}\backslash\{n\}\equiv G_{N}^{(k)}\backslash\{n\}$ where $L^{*}\subseteq[L]$. Then, $G^{k(n)}_{n,N}\cup G^{k(n^{\prime})}_{n^{\prime},N}=G^{(k)}_{N}$ and is independent of $n$. Thus, we can claim $|J_{n,N}\cap J_{n^{\prime},N}|\backslash G_{N}^{(k)}|\geq(L-1)-|G_{N}^{(k)}|\geq\frac{L}{K}-1\geq\ell_{N}$ for $K>1$. Notably, the graph in part 2 of Assumption 1 is connected.
@@
Similarly, we verify that Assumption 1 holds for B-sBMDS under the conditions $B\geq 2\ell_{N}+1$ bands and $B\ll N$. Again, let $\ell_{N}=ceiling(\frac{\sqrt{N}}{2})$, so that it satisfies part 3 of Assumption 1. For all $n\in[N],|J_{n,N}|\geq B$ and for two consecutive objects $(n^{\prime}<n)\in[N],|J_{n^{\prime},N}\cap J_{n,N}|\geq B-1$. Let the number of partitions $K=\frac{\text{\# of observations in the sparsest row}}{\ell_{N}}=\frac{B}{\ell_{N}}$. If $K\notin\mathbb{N}$, then $K$ is the quotient from integer division, and the modulus is distributed evenly among the $\{1,\dots,k-1,k+1,\dots K\}$ remaining partitions. We can now claim that $|G_{n,N}^{(k)}|=\ell_{n}$ and let $G_{n,N}^{(k)}=\{n+1,..,n+\ell_{N}\}$. Thus, $G_{n^{\prime},N}^{(k)}\cup G_{n,N}^{(k)}=\{n,G_{n,N}^{(k)}\}$ and $|J_{n^{\prime},N}\cap J_{n,N}\backslash(G_{n^{\prime},N}^{(k)}\cup G_{n,N}^{(k)})|\geq(B-1)-(|G_{n,N}^{(k)}|)=B-\ell_{N}-1\geq\ell_{N}$ when $B>2\ell_{N}+1$. We obtain a connected graph, fulfilling part 2 of Assumption 1.
@@
Fix $0.10<\alpha<$ 0.1 and $K\in\mathbb{N}$. Let the sequences $\{J_{n,N}\}$, $\{G_{n,N}^{(k)}\}$ and $\{\ell_{N}\}$ satisfy Assumption 1. Let $\epsilon_{N}=\ell_{N}^{-0.5+\alpha}$. Let $(x_{1}^{(N)},\dots,x_{N}^{(N)})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N_{I}(0,1)$ and let $\{\delta_{n,n^{\prime}}^{(N)}\}_{1\leq n^{\prime}<n\leq N}$ be sampled from model (6). Finally, let $(\tilde{x}_{1}^{(N)},\dots,\tilde{x}_{N}^{(N)})\sim p(\cdot|\{\delta_{n,n^{\prime}}^{(N)}\}_{n^{\prime}\in J_{n,N}})$ be sampled from the associated marginal posterior distribution of the model. Then there exists $C>0$ so that the event
@@
See Section A.4 for a short discussion of how similar results may be obtained in fixed dimensions greater than 1.
2.4 Bayesian inference
@@ 3.1 Simulation studies
For a full Bayesian analysis, we put a D-dimensional multivariate normal distribution with mean $\mathbf{0}$ and diagonal covariance matrix $\mathbf{\Lambda}$ as the prior for $\mathbf{x}_{n}$, independently for $n=1,...,N$. The prior for the BMDS error variance $\sigma^{2}$ is an inverse gamma with rate $a$ and shape $b$. One can define hyperpriors for $\mathbf{\Lambda},a,b$, but we assume those parameters are fixed and known in this section. For our simulations, we set $\mathbf{\Lambda}$ equal to the identity $\mathbf{I}_{2},a=1,b=1$ and generate a $N\times 2$ “true” location matrix $\mathbf{X}$ from standard normal distributions. We use $\mathbf{X}$ to calculate the Euclidean distance between pairs $(n,n^{\prime})$ to form the “true” distance matrix, $\bm{\Delta}^{(true)}=\{\delta_{nn^{\prime}}^{(true)}\}$. To create the observed distance matrix $\bm{\Delta}=\{\delta_{nn^{\prime}}\}$, we add i.i.d. noise using a truncated normal distribution with mean 0 and variance $\sigma^{2(true)}$ to the true distance matrix.
@@ 3.1.1 Accuracy
We test the accuracy of the sBMDS models by comparing the simulated “true” distances to those obtained from HMC using the sBMDS posteriors and gradients. Given $S$ iterations, we calculate the mean of the mean squared error ($\overline{\mbox{MSE}}$) as $\overline{\mbox{MSE}}=\frac{1}{S}\sum_{s=1}^{S}\frac{1}{m^{*}}\sum_{n\neq n^{\prime}}(\delta_{nn^{\prime}}^{*(s)}-\delta_{nn^{\prime}}^{(true)})^{2}$ where $\delta_{nn^{\prime}}^{*(s)}$ is the Euclidean distance calculated from the inferred locations of object $n$ and object $n^{\prime}$ at iteration $s$, $\delta_{nn^{\prime}}^{(true)}$ is the “true” distance, and $m^{*}$ is the number of dissimilarities. When the number of objects is greater than 1,000, we randomly sample 1,000 of them to calculate $\overline{\mbox{MSE}}$. We compare distances instead of locations because the locations are not identifiable under distance preserving transformations. We set $\sigma^{(true)}$ to either $0.1,0.2,0.3$ or $0.4$ to change noise levels and run 110,000 iterations, discarding the first 10,000 as burn-in and retaining every 100th iteration. We establish the initial conditions of the latent locations within HMC from classical MDS output.
Figure 1: The mean of the mean squared error (MSE) across all distances using $1$ to $10$ bands for $10$ data points and $1$ to $20$ bands for $100$ and $1{,}000$ data points. We estimate distances from the inferred locations obtained using an adaptive Hamiltonain Monte Carlo algorithm under banded sparse Bayesian multidimensional scaling (B-sBMDS). $\sigma^{2(true)}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma^{(true)}$ corresponds to the BMDS error standard deviation $\sigma$.
Figure 2: The mean of the mean squared error (MSE) from $1{,}000$ randomly sampled distances of distance matrices with $10{,}000$ data points. We estimate distances from the inferred locations obtained using an adaptive Hamiltonian Monte Carlo algorithm under both sparse Bayesian multidimensional scaling (sBMDS) variants, banded sBMDS (B-sBMDS) and landmark sBMDS (L-sBMDS) with $1$ to $20$ bands/landmarks. $\sigma^{2(true)}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma^{(true)}$ corresponds to the BMDS error standard deviation $\sigma$.
Figure 3: Procrustes aligned means of the inferred locations across $100{,}000$ iterations under the B-sBMDS and L-sBMDS frameworks when the number of bands/landmarks is ten and five. We simulate the latent locations from a two-dimensional standard normal distribution and assign a color according to their x-coordinate. $\sigma^{2(true)}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma^{(true)}$ corresponds to the BMDS error standard deviation $\sigma$.
@@ 3.1.2 Computational performance
To better understand the computational benefits of the sBMDS variants, we first calculate the log-likelihood and log-likelihood gradient using B-sBMDS and L-sBMDS for a 10,000 by 10,000 distance matrix. Recall that the number of couplings decreases per additional band/landmark. As a result, we see a parabolic-like relationship between evaluation time (in seconds) and the number of bands/landmarks (Figure 4). If we were to plot the number of couplings vs seconds per evaluation, we would observe linear associations instead. When the number of bands/landmarks is 10,000, we return to the full case. We observe likelihood (gradient) speedups of 457-fold (773-fold), 91-fold (71-fold), 7-fold (10-fold) and 1.3-fold (1.3-fold) for 5, 50, 500 and 5000 bands (landmarks); there appears to be negligible time differences between B-sBMDS and L-sBMDS. Figure 5 emphasizes this correspondence between speedups and number of bands, demonstrating the performance gains using a small number of bands relative to the number of objects. We only scale up to 50 bands because these are reasonable band counts to achieve high accuracy (Figure 1 and 2). We see similar patterns for landmarks in Figure B.2 (Appendix B).
@@
To compare computational performances, we set $\sigma^{(true)}=0.2$, a value that will allow us to establish accurate results while obtaining high acceptance probabilities. We fix the number of bands/landmarks to 10 based on the findings from both Figure 1 and B.1, which confirm that this number ensures high model accuracy when $\sigma^{(true)}=0.2$ and $N<1{,}000$. We then conduct MH and HMC under the full BMDS, B-sBMDS, and L-sBMDS models. For a fair comparison, we run all chains until the minimum effective sample size (ESS) is at least 100. ESS is a function of asymptotic auto-correlation, $\text{ESS}=\frac{S}{1+2\sum_{t=1}^{\infty}\rho_{t}}$, where $\rho_{t}$ is the autocorrelation between samples separated by a lag of $t$ timesteps and $S$ is the length of a time series input. We calculate ESS using the coda package (Plummer et al.,, 2006) in R. We define efficiency as the minimum ESS per hour and take the natural log of it to allow comparison across scales. Figure 6 compares efficiency across the three models and two MCMC algorithms. The sBMDS variants under HMC outperform the others even in moderately high dimensions. MH begins to break down as the number of data points increases because, while it is computationally faster than HMC, the large dimension of the state space prevents efficient exploration, leading to high auto-correlation and low ESS values.

### v3

@@ 2.3 Posterior consistency
$\delta_{nn^{\prime}}\sim N_{(0,M)}(\delta_{nn^{\prime}}^{*},\sigma^{2}),\ n\neq n^{\prime},\ n,n^{\prime}\in[N],$ (6)
a generalization of (1) insofar as M can be any number within the interval $(0,\infty)$. Let the latent locations $\mathbf{X}$ be sampled from a range of values in the interval $I$. The posterior density function of the unknown parameters ($\mathbf{X},I,\sigma^{2},M$) is proportional to $\mathcal{L}(\bm{\Delta}|\mathbf{X},\sigma^{2},M)$, the BMDS likelihood function of model (6), and the priors put on each auxiliary parameter, e.g.,
@@
We examine the posterior consistency of subsampled dissimilarity matrices under simple conditions. Fixing some interval $I$, we sample points $x_{1},\dots,x_{N}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N_{I}(0,1)$. Let $\Delta^{*}$ be the associated Euclidean distance matrix with entries $\delta^{*}_{nn^{\prime}}=|x_{n}-x_{n^{\prime}}|$ and $\delta_{nn^{\prime}}$ be the truncated noisy observations of this matrix sampled from model (6). We set a prior on $I$, $M$ and $\sigma^{2}$ that has compact support and is bounded away from 0 and infinity on its support, e.g., Oh and Raftery, (2001); Man-Suk and Raftery, (2007). In addition, we fix in advance a collection of indices $J_{n,N}\subset[N]\backslash\{n\}$ of observations to keep for each object $n$, treating this choice as non-random in the following. Next, we make some assumptions about which observations are kept.
@@
Fix $K\in\mathbb{N}$. Assume there exists a sequence $\{\ell_{N}\}_{N\in\mathbb{N}}$ and a collection of partitions $\{G_{n,N}^{(k(n))}\}_{k=1}^{K}$ of $J_{n,N}$ with the following properties:
@@
We verify that Assumption 1 holds for L-sBMDS given $L\geq(\ell_{N}+1)K$ landmarks and $L\ll N$. We can think of $\ell_{N}$ as the number of retained entries in the sparsest row (up to a universal constant). Let $\ell_{N}=\lceil\frac{\sqrt{N}}{2}\rceil$, so that it satisfies part 3 of Assumption 1. For all $n\in[N],|J_{n,N}|\geq L$ and for objects $n,n^{\prime}\in[N],|J_{n,N}\cap J_{n^{\prime},N}|\geq L-1$. Consider the partition $G_{n,N}^{(k(n))}=[\lfloor\frac{(k-1)L}{K}\rfloor+1,\lfloor\frac{kL}{K}\rfloor]\backslash\{n\}$. Then for any object $n,n^{\prime}\in[N]$, $G^{(k(n))}_{n,N}\cup G^{(k(n^{\prime}))}_{n^{\prime},N}=\{\lfloor\frac{(k-1)L}{K}\rfloor+1,...,\lfloor\frac{kL}{K}\rfloor\}\equiv G_{N}^{(k)}$ is independent of $n,n^{\prime}$ and of size $\frac{L}{K}$, which satisfies part 1 of Assumption 1. Thus in the most conservative case, e.g., $n\in[L]$ and $n^{\prime}\notin[L]$,
$|J_{n,N}\cap J_{n^{\prime},N}\backslash(G^{(k(n))}_{n,N}\cup G^{(k(n^{\prime}))}_{n^{\prime},N})|=|\{1,...,n-1,n+1,...,L\}\backslash G_{N}^{(k)}|$
$=(L-1)-\frac{L}{K}=\frac{L(K-1)}{K}-1.$
As a result, $|J_{n,N}\cap J_{n^{\prime},N}\backslash(G^{(k(n))}_{n,N}\cup G^{(k(n^{\prime}))}_{n^{\prime},N})|\geq\ell_{N}$ when $L\geq(\ell_{N}+1)K$ for $K>1$. Notably, the graph in part 2 of Assumption 1 is connected.
@@
Similarly, we verify that Assumption 1 holds for B-sBMDS given $B\geq 2\ell_{N}+1$ bands and $B\ll N$. Again, let $\ell_{N}=\lceil\frac{\sqrt{N}}{2}\rceil$, so that it satisfies part 3 of Assumption 1. Under B-sBMDS, the ends of a distance matrix have the fewest indices. To ensure that the size of each partition is at least $\ell_{N}$ at these boundaries, let the number of partitions be $K=\lfloor\frac{B}{\ell_{N}}\rfloor$, satisfying part 1 of Assumption 1. For all $n\in[N],|J_{n,N}|\geq B$ and for two consecutive objects $(n<n^{\prime})\in[N],|J_{n,N}\cap J_{n^{\prime},N}|\geq B-1$. To remove the minimal number of common indices between two consecutive objects, let $G_{n,N}^{(k(n))}=\{n+1,..,n+\ell_{N}\}$ so that $|G_{n,N}^{(k(n))}|=\ell_{N}$. Then, $G_{n,N}^{(k(n))}\cup G_{n^{\prime},N}^{(k(n^{\prime}))}=\{n^{\prime},n^{\prime}+1,...,n^{\prime}+\ell_{N}\}$ and the cardinality of the intersection is $\ell_{N}+1$. Finally,
$|J_{n,N}\cap J_{n^{\prime},N}\backslash(G_{n,N}^{(k(n))}\cup G_{n^{\prime},N}^{(k(n^{\prime}))})|\geq|\{n^{\prime}+1,...,n+B\}\backslash\{n^{\prime},n^{\prime}+1,...,n^{\prime}+\ell_{N}\}|$
|$\geq|\{n^{\prime}+\ell_{N}+1,...,n+B\}|\geq(B-1)-\ell_{N}$
because we remove $\ell_{N}$ elements from the union containing at least $B-1$ elements. Thus when $B\geq 2\ell_{N}+1$, $|J_{n,N}\cap J_{n^{\prime},N}\backslash(G_{n,N}^{(k(n))}\cup G_{n^{\prime},N}^{(k(n^{\prime}))})|\geq\ell_{N}$, and we obtain a connected graph, fulfilling part 2 of Assumption 1.
@@
Fix $0.10<\alpha<$ 0.1 and $K\in\mathbb{N}$. Let the sequences $\{J_{n,N}\}$, $\{G_{n,N}^{(k(n))}\}$ and $\{\ell_{N}\}$ satisfy Assumption 1. Let $\epsilon_{N}=\ell_{N}^{-0.5+\alpha}$. Let $(x_{1}^{(N)},\dots,x_{N}^{(N)})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N_{I}(0,1)$ and let $\{\delta_{nn^{\prime}}^{(N)}\}_{1\leq n<n^{\prime}\leq N}$ be sampled from model (6). Finally, let $(\tilde{x}_{1}^{(N)},\dots,\tilde{x}_{N}^{(N)})\sim p(\cdot|\{\delta_{nn^{\prime}}^{(N)}\}_{n^{\prime}\in J_{n,N}})$ be sampled from the associated marginal posterior distribution of the model. Then there exists $C>0$ so that the event
@@
Theorem 1 proves that we can achieve posterior consistency for latent locations estimated from a subsampled dissimilarity matrix as we are able to recover the estimated latent locations $\tilde{x}_{n}^{(N)}$ up to an additive error of $\mathcal{O}(\epsilon_{N})$ relative to the true latent locations $x_{n}^{(N)}$. We acknowledge that the biggest limitation of this proof is the assumption that we have one-dimensional latent objects. See Section A.4 for a short discussion of how similar results may be obtained in fixed dimensions greater than 1.
2.4 Bayesian computation
@@ 3.1 Simulation studies
For a full Bayesian analysis, we put a D-dimensional multivariate normal distribution with mean $\mathbf{0}$ and diagonal covariance matrix $\mathbf{\Lambda}$ as the prior for $\mathbf{x}_{n}$, independently for $n=1,...,N$. The prior for the BMDS error variance $\sigma^{2}$ is an inverse gamma with rate $a$ and shape $b$. One can define hyperpriors for $\mathbf{\Lambda},a,b$, but we assume those parameters are fixed and known in this section. For our simulations, we set $\mathbf{\Lambda}$ equal to the identity $\mathbf{I}_{2},a=1$ and $b=1$ so that $\mathbf{x}_{n}\sim N(0,\mathbf{I}_{2})$ and $\sigma^{2}\sim IG(1,1)$. To create the observed dissimilarity matrix $\bm{\Delta}=\{\delta_{nn^{\prime}}\}$, we add i.i.d. noise using a truncated normal distribution with mean 0 and variance $\sigma^{2}_{true}$ to a “true” distance matrix. For the “true” distance matrix, $\bm{\Delta}^{(true)}=\{\delta_{nn^{\prime}}^{(true)}\}$, we generate a $N\times 2$ “true” location matrix $\mathbf{X}$ from standard normal distributions and use $\mathbf{X}$ to calculate the Euclidean distance between pairs $(n,n^{\prime})$.
@@ 3.1.1 Accuracy
We test the accuracy of the sBMDS models by comparing the simulated “true” dissimilarities to those obtained from HMC using the sBMDS posteriors and gradients. Given $S$ iterations, we calculate the mean of the mean squared error ($\overline{\mbox{MSE}}$) as $\overline{\mbox{MSE}}=\frac{1}{Sm}\sum_{s=1}^{S}\sum_{n\neq n^{\prime}}(\delta_{nn^{\prime}}^{*(s)}-\delta_{nn^{\prime}}^{(true)})^{2}$ where $\delta_{nn^{\prime}}^{*(s)}$ is the Euclidean distance calculated from the inferred locations of object $n$ and object $n^{\prime}$ at iteration $s$, $\delta_{nn^{\prime}}^{(true)}$ is the “true” Euclidean distance, and $m=N(N-1)/2$ is the number of dissimilarities. We compare distances instead of locations because the locations are not identifiable under distance preserving transformations. For computational convenience, when the number of objects is greater than 1,000, we randomly sample 1,000 distances to calculate $\overline{\mbox{MSE}}$. We set $\sigma_{true}$ to either $0.1,0.2,0.3$ or $0.4$ to change noise levels and run 110,000 iterations, discarding the first 10,000 as burn-in and retaining every 100th iteration. We establish the initial conditions of the latent locations within HMC from classical MDS output.
Figure 1: The mean of the mean squared error (MSE) across all distances using $1$ to $10$ bands for $10$ data points and $1$ to $20$ bands for $100$ and $1{,}000$ data points. We estimate Euclidean distances from the inferred locations obtained using an adaptive Hamiltonian Monte Carlo algorithm under banded sparse Bayesian multidimensional scaling (B-sBMDS). $\sigma^{2}_{true}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma_{true}$ corresponds to the BMDS error standard deviation $\sigma$.
Figure 2: The mean of the mean squared error (MSE) across $1{,}000$ distances, randomly sampled from distance matrices with $10{,}000$ data points. We estimate Euclidean distances from the inferred locations obtained using an adaptive Hamiltonian Monte Carlo algorithm under both sparse Bayesian multidimensional scaling (sBMDS) variants, banded sBMDS (B-sBMDS) and landmark sBMDS (L-sBMDS) with $1$ to $20$ bands/landmarks. $\sigma^{2}_{true}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma_{true}$ corresponds to the BMDS error standard deviation $\sigma$.
Figure 3: Procrustes aligned means of the inferred locations across $100{,}000$ iterations under the B-sBMDS and L-sBMDS frameworks when the number of bands/landmarks is ten and five. The number of data points is 1,000. We simulate the latent locations from a two-dimensional standard normal distribution and assign a color according to their x-coordinate. $\sigma^{2}_{true}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma_{true}$ corresponds to the BMDS error standard deviation $\sigma$.
@@
3.1.2 Sensitivity to model misspecification
To observe how the sparse variants behave under model misspecifications, we explore two possible situations: 1) a mismatch between the true dimensionality and that specified by the scientist and 2) heavy-tailed, rather than truncated normal, noise. For case 1, we vary the dimension of the “true” location matrix from 2 to 10 while fixing the embedding dimension to 2 and the truncated Gaussian noise variance $\sigma^{2}_{true}$ to 0.2. As expected, $\overline{\mbox{MSE}}$ decreases as the true underlying dimensionality approaches the embedding dimension. Full BMDS and B-sBMDS with 20 bands are more robust to dimension misspecification than classical MDS, and the accuracy for B-sBMDS with 20 bands closely matches that of full BMDS (Figure 4).
For case 2, we assume a correctly-specified-dimensional Euclidean space, but add i.i.d log-normal noise to the “true” distance matrix. We bootstrap the $\overline{\mbox{MSE}}$ across all distances from 100 data points, a 100 times and plot the mean of $\overline{\mbox{MSE}}$ along with error bars representing $\pm$ the standard deviation of $\overline{\mbox{MSE}}$. Figure 5 demonstrates that, even with heavy-tailed data, both sparse variants achieve comparable $\overline{\mbox{MSE}}$s to full BMDS’s at a low number of bands/landmarks. In both cases, we observe that B-sBMDS seems to be less sensitive to model misspecification than L-sBMDS.
Figure 4: The mean of the mean squared error (MSE) across all distances using $20$ bands/landmarks for 100 and 1,000 data points. We vary the dimension space of the “true” latent locations while fixing the latent dimensionality to two. We estimate Euclidean distances from the inferred locations obtained using an adaptive Hamiltonian Monte Carlo algorithm under the B-sBMDS, L-sBMDS and full BMDS frameworks. Additionally, we compare the mean MSE across all distances from the inferred locations using classical MDS.
Figure 5: The average mean of the mean squared error (MSE) across all distances from 100 data points evaluated at intervals of 10 bands/landmarks, from 10 to 100, repeated 100 times. The dot is the average mean of MSE, and the error bars are $\pm$ one standard deviation away from this mean. We estimate Euclidean distances from the inferred locations obtained using an adaptive Hamiltonian Monte Carlo algorithm under both sparse Bayesian multidimensional scaling (sBMDS) variants, banded sBMDS (B-sBMDS) and landmark sBMDS (L-sBMDS). When the number of bands (landmarks) equals 100, we return to full BMDS. $\sigma^{2}_{true}$ is the variance component of the log-normal noise centered at 00 added to the “true” distance matrix such that the distribution of the observed distance matrix has heavy-tails.
@@ 3.1.3 Computational performance
To better understand the computational benefits of the sBMDS variants, we first calculate the log-likelihood and log-likelihood gradient using B-sBMDS and L-sBMDS for a 10,000 by 10,000 Euclidean distance matrix. Recall that the number of couplings decreases per additional band/landmark. As a result, we see a parabolic-like relationship between evaluation time (in seconds) and the number of bands/landmarks (Figure 6). If we were to plot the number of couplings vs seconds per evaluation, we would observe linear associations instead. When the number of bands/landmarks is 10,000, we return to the full case. We observe likelihood (gradient) speedups of 457-fold (773-fold), 91-fold (71-fold), 7-fold (10-fold) and 1.3-fold (1.3-fold) for 5, 50, 500 and 5000 bands (landmarks); there appears to be negligible time differences between B-sBMDS and L-sBMDS. Figure 7 emphasizes this correspondence between speedups and number of bands, demonstrating the performance gains using a small number of bands relative to the number of objects. We observe approximately 3-fold, 10-fold and 40-fold speedups when applying the sBMDS likelihoods and gradients to 500, 1,000 and 5,000 data points with 50 bands/landmarks. We only scale up to 50 bands because these are reasonable band counts to achieve high accuracy (Figure 1 and 2). We see similar patterns for landmarks in Figure B.2 (Appendix B).
@@
To compare computational performances, we set $\sigma_{true}=0.2$, a value that will allow us to establish accurate results while obtaining high acceptance probabilities. We fix the number of bands/landmarks to 10 based on the findings from both Figure 1 and B.1, which confirm that this number ensures high model accuracy when $\sigma_{true}=0.2$ and $N<1{,}000$. We then conduct MH and HMC under the full BMDS, B-sBMDS, and L-sBMDS models. For a fair comparison, we run all chains until the minimum effective sample size (ESS) is at least 100. ESS is a function of asymptotic auto-correlation, $\text{ESS}=\frac{S}{1+2\sum_{t=1}^{\infty}\rho_{t}}$, where $\rho_{t}$ is the autocorrelation between samples separated by a lag of $t$ timesteps and $S$ is the length of a time series input. We calculate ESS using the coda package (Plummer et al.,, 2006) in R. We define efficiency as the minimum ESS per hour and take the natural log of it to allow comparison across scales. Figure 8 compares efficiency across the three models and two MCMC algorithms. The sBMDS variants under HMC outperform the others even in moderately high dimensions. MH begins to break down as the number of data points increases because, while it is computationally faster than HMC, the large dimension of the state space prevents efficient exploration, leading to high auto-correlation and low ESS values.